When properly detonated, ammonium nitrate explodes violently, releasing hot gases: NH4NO3(s)->N2O(g)+2 H2O. Calculate the total volume of gas released at 975 degree C and 1.22 atm by the explosion of 46.4 grams of ammonium nitrate.
1
Expert's answer
2018-01-29T13:58:07-0500
NH₄NO₃=N₂O+2H₂O
m(N₂O)=m(NH₄NO₃)*M(N₂O)/M(NH₄NO₃)=24.9g
V(N₂O)=24.9*22.4/44=12.67L
V=101.325*12.67*(975+273)/123616.5*273=57.48L
1.22atm=123616.5Pa